# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series

### v1

@@ Abstract
We analyze a high-frequency, real-world compositional dataset: Airbnb’s daily currency-fee proportions across multiple regions. These proportions sum to one and exhibit changing levels of volatility, particularly following disruptions such as the COVID-19 pandemic. Standard compositional models often assume constant variance, but our data reveal episodes of elevated uncertainty and clustering in volatility.
Motivated by these empirical patterns, we propose a Bayesian Dirichlet Auto-Regressive Moving Average framework extended with a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component—the B-DARMA-DARCH model. This novel approach retains the Dirichlet likelihood, ensuring valid predictions on the simplex, and introduces a time-varying precision parameter via an ARMA recursion. As a result, B-DARMA-DARCH captures both mean dynamics and stochastic volatility in compositional time series.
Simulations and empirical tests on the Airbnb currency data show that B-DARMA-DARCH outperforms simpler Dirichlet ARMA and Bayesian log-ratio VARMA models in forecasting accuracy, residual diagnostics, and interval coverage. By accommodating volatility clustering and structural breaks, B-DARMA-DARCH provides a robust framework for compositional data analysis in finance and beyond, supporting improved risk management and uncertainty quantification in settings where proportions and their volatility are of primary interest.
@@ 1 Introduction
In contrast, the Dirichlet distribution models compositional data directly on the simplex. Benjamin et al. (2003) proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al. (1993), da Silva et al. (2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al. (2024, 2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
@@
Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH). Unlike existing Dirichlet ARMA methods that assume a fixed or purely deterministic precision parameter, our framework introduces a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component. This addition endows the standard Bayesian Dirichlet ARMA model with a GARCH-like recursion on the precision $\phi_{t}$, enabling it to:
@@
    Model volatility clustering by letting $\phi_{t}$ depend stochastically on recent innovations and prior volatility, thereby reflecting abrupt regime changes and prolonged uncertainty.
@@
These innovations allow B-DARMA-DARCH to simultaneously handle both systematic and stochastic changes in compositional data, providing a flexible, unified approach to dynamic heteroskedasticity on the simplex—a capability hitherto missing in Bayesian compositional time-series models.
We apply the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, demonstrating its effectiveness in capturing complex temporal dynamics and volatility patterns in real-world compositional data. We conduct simulation studies comparing the B-DARMA-DARCH model with the standard B-DARMA and Bayesian transformed VARMA (B-tVARMA) models. The B-DARMA-DARCH model consistently outperforms the other models in forecast accuracy and residual diagnostics, highlighting its robustness and applicability in scenarios with heteroskedasticity, structural breaks, and regime changes. By directly modeling compositional data within the simplex and incorporating time-varying volatility, our proposed B-DARMA-DARCH model offers a powerful tool for analyzing dynamic compositional time series.
The next section presents the B-DARMA-DARCH model. Section 3 presents simulation studies comparing the B-DARMA-DARCH to both the B-DARMA data model and a Bayesian transformed data normal VARMA model. Section 4 presents analysis of the Airbnb data. The paper closes with a short discussion.
@@ 2.1 DARMA Data model
$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t}\sim\operatorname{Dirichlet}\left(\phi_{t}\boldsymbol{\mu}_{t}\right),$ (1)
@@ 2.1.1 Modeling the Mean using the Additive Log-Ratio Transformation
where $\boldsymbol{\eta}_{t}$ is the linear predictor, a $J-1$-vector taking values in $\mathbb{R}^{J-1}$, $j^{*}$ is a chosen reference component $1\leq j^{*}\leq J$, and the element of $\boldsymbol{\eta}_{t}$ that would correspond to the $j^{*}$th element $\log(\mu_{j^{*}}/\mu_{j^{*}})=0$ is omitted. Given $\boldsymbol{\eta}_{t}$, $\boldsymbol{\mu}_{t}$ is defined by the inverse of equation (3) where $\mu_{tj}=\exp(\eta_{tj})/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$ for $j=1,\dots,J$, $j\neq j^{*}$ and for $j=j*$, $\mu_{tj*}=1/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$.
The alr-transformed mean vector $\mathbf{\eta}_{t}$ is modeled as a Vector Auto-Regressive Moving Average (VARMA) process
@@ 2.1.2 Precision Parameter $\phi_{t}$
In the B-DARMA model, precision parameter $\phi_{t}$ is modeled with log link as a function of an $r_{\gamma}$-vector of covariates $\boldsymbol{z}_{t}$,
$\phi_{t}=\exp(\boldsymbol{z}_{t}\boldsymbol{\gamma}),$ (5)
@@ 2.2 DARCH Process for Precision Parameter $\phi_{t}$
For ease of notation, we set $P=L$ and $Q=K$, but this assumption is easily relaxed as shown in section 3. The DARCH component models the volatility in the precision parameter $\phi_{t}$, addressing non-constant variability over time. By capturing volatility clustering and external influences, the DARCH process enhances the model’s ability to understand compositional data dynamics, improving forecasting and risk management.
Alternative log-ratio transformations, such as the centered log-ratio (CLR) or the isometric log-ratio (ILR), can also be used. A detailed discussion of these transformations is provided in the Supplementary Material (Section Alternative Log-Ratio Transformations).
@@ 2.3 Joint Predictive Distribution
Define the $C$-vector $\boldsymbol{\theta}$ of all unknown parameters $\boldsymbol{\theta}=(\boldsymbol{A}_{prs},\boldsymbol{B}_{qrs},\boldsymbol{\beta}^{\prime},\boldsymbol{\gamma}^{\prime},\tau_{j},\alpha_{i})^{\prime}$, where $r,s=1,\ldots,J-1$ index all elements of matrices $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$, $p=1,\ldots,P$, $q=1,\ldots,Q$, $i=1,\dots,L$, $j=1,\dots,K$ and $C=(P+Q)*(J-1)^{2}+L+K+r_{\beta}+r_{\gamma}$. Our prior beliefs about $\boldsymbol{\theta}$ ($p(\boldsymbol{\theta})$) are updated by Bayes’ theorem to give the posterior
@@
[…]
@@ 2.4 Differences between B-DARMA, B-tVARMA, and B-DARCH
The differences among the Bayesian-ALR-transformed VARMA (B-tVARMA), B-DARMA, and B-DARCH models lie in their handling of compositional constraints and volatility. The B-tVARMA model transforms compositional data to an unconstrained space, which may introduce distortions if model assumptions like normality are violated. In contrast, the B-DARMA and B-DARCH models directly model data within the simplex using the Dirichlet distribution. For volatility, the B-tVARMA model assumes constant variance unless extended to include heteroskedasticity. The B-DARMA model allows a time-varying precision parameter $\phi_{t}$ modeled by exogenous factors but lacks a stochastic structure, potentially missing stochastic volatility patterns. The B-DARCH model introduces a dynamic $\phi_{t}$ with an ARMA structure, capturing both systematic and stochastic changes in volatility, such as clustering and sudden shifts. While the B-tVARMA model is simpler, it may inadequately represent compositional data and volatility dynamics. The B-DARMA model offers a balance but may miss important volatility features. The B-DARCH model, though more complex, comprehensively accounts for mean dynamics and volatility patterns, enhancing model fit and forecasting accuracy when heteroskedasticity is present.
@@ 3.1 Overview
We conducted six simulation studies to evaluate the performance of the B-DARMA(1,0)-DARCH(1,1) model compared to the B-DARMA(1,0) and B-tVARMA(1,0) models. Each study generates data from either a DARMA, DARMA-DARCH, or tVARMA and introduces random shocks to simulate misreported observations (1-3) or redraws values for all parameters (4-6) to simulate a regime shift. Each study simulates 50 datasets, and evaluates out-of-sample forecast accuracy and in-sample residual autocorrelation.
3.2 Simulation Studies Setup
In all six studies, datasets with five compositional components were simulated over $T=100$ time points. The first 60 observations were used for training, and the remaining 40 for testing. For each simulation, a new set of parameters was drawn, giving different intercept vectors, auto-regressive matrices, and precision parameters across simulations.
All six simulation studies have the same underlying mean vector formulation for the latent ALR-transformed process
$\boldsymbol{\eta}_{t}=\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}).$
The auto-regressive coefficient matrix $\boldsymbol{A}$ was a $4\times 4$ matrix with elements drawn from a uniform distribution, $A_{ij}\sim\mathcal{U}(-0.75,0.75)$, independently for each simulation. For each of the 50 simulations, the elements of the intercept vector $\boldsymbol{\beta}$ were generated by drawing five elements $\beta_{j}^{*},j=1,\dots,5$ from $\mathcal{N}(0.2,0.03^{2})$, then normalized to sum to 1 ($\beta_{j}^{*}=\frac{\beta_{j}^{*}}{\sum_{j^{*}=1}^{5}\beta_{j}^{*}})$, transformed using the ALR transformation to give the $4\times 1$ intercept vector $\boldsymbol{\beta}=\text{alr}(\boldsymbol{\beta}^{*})$. For each simulation, the initial values were
$y_{1j}=.2+\epsilon_{j},\ \epsilon_{j}\sim N(0,.01^{2}),\ j=1,\dots,5$
followed by $y_{1}=\frac{y_{1}}{\sum y_{1}}.$
3.2.1 Simulation Studies 1 & 4: DARMA DGP
In Simulations 1 and 4, the data-generating process (DGP) was a DARMA model, $\mathbf{y}_{t}|\mathbf{\mu}_{t},\phi_{t}\sim\text{Dirichlet}(\phi_{t}\mathbf{\mu}_{t})$ with a constant precision parameter $\log(\phi_{t})=\phi_{0}$. The precision parameter $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$
3.2.2 Simulation Studies 2 & 5: DARCH DGP
The data in the second and fifth simulations were generated using a DARMA-DARCH process, where $\phi_{t}$ was generated
$\log(\phi_{t})=\phi_{0}+\alpha(\log(\phi_{t-1})-\phi_{0})+\tau\sum_{i=1}^{4}(\text{alr}(y_{t-1,i})-\eta_{t-1,i})^{2}.$
The true intercept $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$ We fixed parameters $\alpha=0.8$ and $\tau=-0.95$ to simulate data with high volatility persistence and strong responsiveness to past innovations, creating heteroskedasticity patterns that test the B-DARCH model’s ability to capture dynamic volatility.
3.2.3 Simulation Studies 3 & 6: tVARMA DGP
The data in the third and sixth simulations were generated using a tVARMA process with

### v2

# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares

@@ Abstract
In marketplace finance, the daily mix of billing currencies is compositional data that drives forecasting, reporting, and treasury risk. We study Airbnb’s 2017–2020 currency‑fee shares across four regions and present a Bayesian Dirichlet ARMA model with a Dirichlet ARCH volatility component, B‑DARMA‑DARCH. The model keeps predictions on the simplex, models mean dynamics with VARMA terms on the additive log‑ratio scale, and lets precision vary over time so volatility can spike during disruptions and settle as conditions normalize. We include weekly and yearly seasonality and a trend, and we evaluate against Dirichlet ARMA and transformed‑Gaussian VARMA across simulation designs with misreported observations and temporary regime shifts, then validate on a held‑out window and test in 2020 Q4. Across simulations and the application, B‑DARMA‑DARCH delivers the most accurate aggregate forecasts, better‑calibrated intervals, and weaker residual persistence at medium lags. Modeling precision as a dynamic process provides a practical, interpretable way to forecast proportions and quantify uncertainty when the noise itself moves.
@@ 1 Introduction
Let $Q_{t,c}$ denote gross booking amount in currency $c$ on day $t$, and let $\mathrm{FX}_{t,c\rightarrow USD}$ be the USD price of one unit of $c$. Consolidated revenue in USD can be written as
$R_{t}^{USD}\;=\;\sum_{c=1}^{J}Q_{t,c}\,\mathrm{FX}_{t,c\rightarrow USD}\;=\;V_{t}\sum_{c=1}^{J}S_{t,c}\,\mathrm{FX}_{t,c\rightarrow USD},\quad S_{t,c}=\frac{Q_{t,c}}{V_{t}},\;V_{t}=\sum_{c=1}^{J}Q_{t,c}.$ (1)
Thus a USD forecast requires both a scale forecast $V_{t}$ (total nominal volume) and a mix forecast $S_{t,\cdot}$ (currency shares). One can pursue a bottom‑up route (model each $Q_{t,c}$ and normalize to obtain $S_{t,\cdot}$) or a top‑down route (model $S_{t,\cdot}$ directly on the simplex). Our focus is the latter, because the decision‑relevant object for reporting and hedging is the composition itself: even small swings in $S_{t,\cdot}$ materially change $R_{t}^{USD}$ when multiplied by large $V_{t}$ and moving FX rates. Empirically, the share series exhibit recurrent bursts of dispersion (see Figure 1), which motivates an observation‑driven volatility mechanism on the simplex.
In contrast, the Dirichlet distribution models compositional data directly on the simplex. Benjamin et al. (2003) proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al. (1993), da Silva et al. (2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al. (2024, 2025); Katz and Maierhofer (2025); Katz (2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
@@
Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARCH). Unlike existing Dirichlet ARMA methods that assume a fixed or purely deterministic precision parameter, our framework introduces a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component. This addition endows the standard Bayesian Dirichlet ARMA model with a GARCH-like recursion on the precision $\phi_{t}$, enabling it to:
@@
    Model volatility clustering by letting $\phi_{t}$ depend dynamically on recent innovations and prior volatility, thereby reflecting abrupt regime changes and prolonged uncertainty.
@@
These innovations allow B-DARCH to simultaneously handle both systematic and dynamic changes in compositional data, providing a flexible, unified approach to dynamic heteroskedasticity on the simplex—a capability hitherto missing in Bayesian compositional time-series models.
We apply the B-DARCH model to Airbnb’s currency fee proportions across different regions, demonstrating its effectiveness in capturing complex temporal dynamics and volatility patterns in real-world compositional data. We conduct simulation studies comparing the B-DARCH model with the standard B-DARMA and Bayesian transformed VARMA (B-tVARMA) models. The B-DARCH model consistently outperforms the other models in forecast accuracy and residual diagnostics, highlighting its robustness and applicability in scenarios with heteroskedasticity, structural breaks, and regime changes. By directly modeling compositional data within the simplex and incorporating time-varying volatility, our proposed B-DARCH model offers a powerful tool for analyzing dynamic compositional time series.
The remainder of the paper is organized as follows. The next section introduces the B-DARCH model. Section 3 reports simulation studies comparing B-DARCH with three alternatives: (i) a B-DARMA data model; (ii) a Bayesian transformed-data normal VARMA model; and (iii) a time-varying-parameter comparator in ALR space (B-TVP-tVARMA). Section 4 analyzes the Airbnb data. The paper concludes with a brief discussion.
@@ 2.1 DARMA Data model
$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t}\sim\operatorname{Dirichlet}\left(\phi_{t},\boldsymbol{\mu}_{t}\right),$ (2)
@@ 2.1.1 Modeling the Mean using the Additive Log-Ratio Transformation
where $\boldsymbol{\eta}_{t}$ is the linear predictor, a J1J-1-vector taking values in $\mathbb{R}^{J-1}$, $j^{*}$ is a chosen reference component $1\leq j^{*}\leq J$, and the element of $\boldsymbol{\eta}_{t}$ that would correspond to the $j^{*}$th element $\log(\mu_{j^{*}}/\mu_{j^{*}})=0$ is omitted. Given $\boldsymbol{\eta}_{t}$, $\boldsymbol{\mu}_{t}$ is defined by the inverse of equation (4) where $\mu_{tj}=\exp(\eta_{tj})/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$ for $1,\dots,J, jj$$\neq j^{*}$ and for j=jj=j*, $\mu_{tj*}=1/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$.
The ALR-transformed mean vector $\boldsymbol{\eta}_{t}$ is modeled with an observation-driven VARMA-style recursion that uses autoregressive and moving-average terms built from past ${\textnormal{alr}}(\mathbf{y})$ and past linear predictors:
@@
Interpretation.
Equation (5) defines $\boldsymbol{\eta}_{t}$ deterministically conditional on past data and past linear predictors, without introducing a separate innovation term for $\boldsymbol{\eta}_{t}$. Hence it is an observation-driven specification in the sense of GARCH-type recursions rather than a parameter-driven state-space evolution. A natural alternative would be to endow $\boldsymbol{\eta}_{t}$ with a stochastic transition equation (for example, a linear or nonlinear state-space model with evolution noise) and integrate over the latent path. In our Dirichlet likelihood with the DARCH precision recursion, this approach becomes computationally onerous due to the nonlinearity of the simplex mapping and the volatility recursion. The observation-driven form in (5) keeps inference tractable while still allowing rich short-run dynamics through the AR terms in ${\textnormal{alr}}(\mathbf{y})$ and the MA terms in $\boldsymbol{\eta}$. We complement this with a transformed-Gaussian comparator that does include time-variation on the mean side via time-varying coefficients (B-TVP-tVARMA), thereby illustrating the trade-off between state evolution and observation-driven recursions within our framework.
@@ 2.1.2 Precision Parameter $\phi_{t}$
In the B-DARMA model, precision parameter $\phi_{t}$ is modeled with log link as a function of an $r_{\gamma}$-vector of covariates $\mathbf{z}_{t}$,
$\phi_{t}=\exp(\mathbf{z}_{t}\boldsymbol{\gamma}),$ (6)
@@ 2.2 DARCH Process for Precision Parameter $\phi_{t}$
For compactness, we write B-DARCH(P,Q)(P,Q) to denote a B-DARMA mean model of order $(P,Q)$ coupled with a DARCH precision process for $\phi_{t}$ of order $(L,K)$. In general, $(L,K)$ need not equal $(P,Q)$; when they differ, we state both explicitly as B-DARMA(P,Q)(P,Q)–DARCH(L,K)(L,K) (e.g., B-DARMA(1,0)(1,0)–DARCH(1,1)(1,1)). The DARCH component governs time-variation in the precision $\phi_{t}$, accommodating volatility clustering and covariate effects; see Section 3 for cases with $(L,K)\neq(P,Q)$.
Alternative log-ratio transformations, such as the centered log-ratio (CLR) or the isometric log-ratio (ILR), can also be used. A detailed discussion of these transformations is provided in the Supplementary Material (Section 6).
@@ 2.3 Joint Predictive Distribution
Define the $C$-vector $\boldsymbol{\theta}$ of all unknown parameters $\boldsymbol{\theta}=(\mathbf{A}_{prs},\mathbf{B}_{qrs},\boldsymbol{\beta}^{\prime},\boldsymbol{\gamma}^{\prime},\tau_{j},\alpha_{i})^{\prime}$, where $r,s=1,\ldots,J-1$ index all elements of matrices $\mathbf{A}_{p}$ and $\mathbf{B}_{q}$, $p=1,\ldots,P$, $q=1,\ldots,Q$, $i=1,\dots,L$, $j=1,\dots,K$ and $C=(P+Q)*(J-1)^{2}+L+K+r_{\beta}+r_{\gamma}$. Our prior beliefs about $\boldsymbol{\theta}$ ($p(\boldsymbol{\theta})$) are updated by Bayes’ theorem to give the posterior
@@
where $p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})=\prod_{t=m+1}^{T}p(\mathbf{y}_{t}|\boldsymbol{\theta},\mathbf{y}_{(t-m):(t-1)})$, $p(\mathbf{y}_{t}|\boldsymbol{\theta},\mathbf{y}_{(t-m):(t-1)})$ is the density of the Dirichlet in (3), and the normalizing constant $p(\mathbf{y}_{(m+1):T}|\mathbf{y}_{1:m})=\int p(\boldsymbol{\theta})p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})d\boldsymbol{\theta}$. Our objective is to forecast the upcoming $S$ observations, denoted as $\mathbf{y}_{(T+1):(T+S)}$. The joint predictive distribution for these future observations is
$p(\mathbf{y}_{(T+1):(T+S)}\mid\mathbf{y}_{1:T})=\int_{\boldsymbol{\theta}}p(\mathbf{y}_{(T+1):(T+S)}\mid\boldsymbol{\theta})\,p(\boldsymbol{\theta}\mid\mathbf{y}_{1:T})\,d\boldsymbol{\theta}.$
@@ 2.4 Differences between B-DARMA, B-tVARMA, and B-DARCH
The three models differ in how they respect the simplex and how they treat time‑varying volatility. B‑tVARMA moves the composition to the ALR space and fits a Gaussian VARMA with constant observation covariance unless explicitly extended. This keeps estimation simple but can misalign with the simplex after back‑transformation if the Gaussian–ALR assumptions are strained. B‑DARMA keeps the Dirichlet likelihood on the simplex; its precision $\phi_{t}$ is a deterministic regression in covariates, so dispersion may change with seasonality or trend but does not react to forecast errors. B‑DARCH also retains the Dirichlet likelihood but replaces the fixed precision with a GARCH‑type, observation‑driven recursion for log $\phi_{t}$. Volatility is updated dynamically from past residuals and past precision, allowing clustering and short‑run bursts in dispersion while preserving the compositional support. This is time‑varying volatility in an observation‑driven sense, not a parameter‑driven stochastic‑volatility state‑space model. In practice, B‑tVARMA is parsimonious but may under‑represent compositional and volatility features; B‑DARMA captures the simplex with covariate‑driven variance; and B‑DARCH adds an adaptive volatility mechanism that improves fit and forecasting when conditional heteroskedasticity matters.
Assumption sensitivity and alternatives.
Both B-tVARMA and the Dirichlet-based B-DARMA/B-DARCH are parametric choices that can distort if their distributional assumptions are violated. On the ALR scale, B-tVARMA assumes Gaussian residuals; heavy tails or skewness can lead to bias on back-transformation to the simplex. For B-DARMA and B-DARCH, the Dirichlet likelihood imposes a specific mean–variance relationship and a restricted covariance structure ($\operatorname{Cov}(y_{i},y_{j})=-\mu_{i}\mu_{j}/(\phi_{t}+1)$ for $i\neq j$), which rules out positive contemporaneous correlations among components. When the empirical dispersion or dependence deviates from these constraints, Dirichlet-based models can also misrepresent uncertainty or cross-component co-movement.
Other families for proportions are available, including logistic-normal models on a log-ratio scale, generalized Dirichlet specifications, mixtures (e.g., Dirichlet mixtures), and zero-augmented variants for sparse compositions. We focus on the Dirichlet class because it guarantees simplex-valid predictions and enables a tractable, observation-driven volatility mechanism via the DARCH recursion. The transformed-Gaussian comparator (B-tVARMA) provides a complementary baseline with a richer covariance structure in ALR space. In applications where positive cross-correlations or more flexible dispersion are essential, generalized Dirichlet or logistic-normal formulations are natural extensions that can be substituted into the same modeling workflow.
@@ 3.1 Overview
We evaluate forecasting performance under two generative mechanisms for compositional time series—Dirichlet ARMA (DARMA) and transformed‑data Gaussian VARMA (tVARMA)—each under two perturbations (misreported observations; temporary regime shifts), yielding four simulation studies in total. In addition to B-DARMA(1,0), B-tVARMA(1,0), and the volatility‑aware B-DARCH(1,0), we also consider a time‑varying‑parameter comparator in ALR space (B-TVP-tVARMA) with
$\boldsymbol{\eta}_{t}=\sum_{p=1}^{P}\mathbf{A}_{t,p}\!\big(\mathbf{Y}_{t-p}-\mathbf{X}_{t-p}\boldsymbol{\beta}\big)+\sum_{q=1}^{Q}\mathbf{B}_{t,q}\!\big(\mathbf{Y}_{t-q}-\boldsymbol{\eta}_{t-q}\big)+\mathbf{X}_{t}\boldsymbol{\beta},\qquad\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\boldsymbol{\Sigma}),$
and constant observation covariance $\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma})$. The time‑varying coefficient matrices follow mean‑reverting AR(1) evolutions around long‑run targets,
$\mathbf{A}_{t,p}=\bar{\mathbf{A}}_{p}+\rho^{(A)}_{p}\!\big(\mathbf{A}_{t-1,p}-\bar{\mathbf{A}}_{p}\big)+\tau^{(A)}_{p}\,\boldsymbol{Z}^{(A)}_{t,p},\qquad\mathbf{B}_{t,q}=\bar{\mathbf{B}}_{q}+\rho^{(B)}_{q}\!\big(\mathbf{B}_{t-1,q}-\bar{\mathbf{B}}_{q}\big)+\tau^{(B)}_{q}\,\boldsymbol{Z}^{(B)}_{t,q}$,
with $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t,p})$ and $\operatorname{vec}(\boldsymbol{Z}^{(B)}_{t,q})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},\mathbf{I})$. For the simulation studies we use the B-TVP-VAR(1) special case (P=1P{=}1, Q=0Q{=}0) with
$\boldsymbol{\eta}_{t}=\mathbf{A}_{t,1}\!\big(\mathbf{Y}_{t-1}-\mathbf{X}_{t-1}\boldsymbol{\beta}\big)+\mathbf{X}_{t}\boldsymbol{\beta},\qquad\mathbf{A}_{t,1}=\bar{\mathbf{A}}_{1}+\rho^{(A)}_{1}\!\big(\mathbf{A}_{t-1,1}-\bar{\mathbf{A}}_{1}\big)+\tau^{(A)}\,\boldsymbol{Z}^{(A)}_{t}$,
where $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\tau^{(A)}\sim\mathrm{Half\!-\!Normal}(0,\,0.1^{2})$.
Across all studies we report out‑of‑sample forecasting accuracy and in‑sample residual autocorrelation.
3.2 Simulation Setup
Each dataset contains five compositional components observed over $T=100$ time points; the first 60 observations form the training set and the remaining 40 the test set. For each simulation, parameters are redrawn (intercepts, autoregressive matrices, and—where applicable—precision or covariance) to ensure heterogeneity across runs. Initialization uses a perturbed balanced composition,
$y_{1j}=0.2+\epsilon_{j},\quad\epsilon_{j}\sim\mathcal{N}(0,\,0.01^{2}),\ j=1,\dots,5,\qquad y_{1}\leftarrow\frac{y_{1}}{\sum_{j=1}^{5}y_{1j}}.$
Across all studies, the latent ALR-scale mean follows
$\boldsymbol{\eta}_{t}\;=\;\boldsymbol{\beta}\;+\;\mathbf{A}\,\big(\mathrm{alr}(\mathbf{Y}_{t-1})-\boldsymbol{\beta}\big),$
with a $4\times 4$ matrix $\mathbf{A}$ drawn elementwise as $A_{ij}\sim\mathcal{U}(-0.75,\,0.75)$ for each simulation. The simplex intercept is obtained by drawing $\beta_{j}^{\ast}\sim\mathcal{N}(0.2,\,0.03^{2})$ for $j=1,\dots,5$, normalizing to $\sum_{j}\beta_{j}^{\ast}=1$, and applying ALR: $\boldsymbol{\beta}=\mathrm{alr}(\boldsymbol{\beta}^{\ast})$.
Comparator defined: B-TVP–VAR(1).
To allow time‑variation in linear dynamics under constant observation variance, the B‑TVP–VAR(1) comparator specifies